addle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations. Standard Custom Direct labor costs $45,000 $113,000 Machine hours 1,140 1,110 Setup hours 106 430 Total estimated overhead costs are $308,000. Overhead cost allocated to the machining activity cost pool is $195,000, and $113,000 is allocated to the machine setup activity cost pool. Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.) Predetermined overhead rate enter the predetermined overhead rate in percentages of direct labor cost rounded to 2 decimal places % of direct labor cost

Compute the overhead rates using the activity-based costing approach. (Round answers to 2 decimal places, e.g. 12.25.)

Machining

$enter a dollar amount per machine hour rounded to 2 decimal places

per machine hour
Machine setup

$enter a dollar amount per setup hour rounded to 2 decimal places

per setup hour

eTextbook and Media

  

  

Determine the difference in allocation between the two approaches. (Round answers to 0 decimal places, e.g. 1,225.)

Traditional costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

Activity-based costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places
Answer #1
1
Total Overhead cost 308000
Divide by Total Direct labor costs 158000 =45000+113000
Predetermined overhead rate 194.94% % of direct labor cost
2
Machining 86.67 per machine hour =195000/(1140+1110)
Machine setup 210.82 per setup hour =113000/(106+430)
3
Traditional costing:
Standard 87723 =45000*194.94%
Custom 220282 =113000*194.94%
Activity-based costing:
Standard 121151 =(1140*86.67)+(106*210.82)
Custom 186856 or 186857 =(1110*86.67)+(430*210.82)